Find the EAR in each of the following cases (Use 365 days a year. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.):


Stated Rate (APR) Number of Times Compounded Effective Rate (EAR)
9.4 % Quarterly %
18.4 Monthly     
14.4 Daily     
11.4 Infinite     
Answer #1

EAR = [1 + (APR / m)]m − 1

EAR = [1 + (0.094 / 4)]4 − 1 = 0.0974 or 9.74%

EAR = [1 + (0.184 / 12)]12 − 1 = 0.2003 or 20.03%

EAR = [1 + (0.144 / 365)]365 − 1 = 0.1549 or 15.49%

EAR = eq − 1 = e0.114− 1 =  0.1208 or 12.08%